# How Real Are Synthetic Therapy Conversations? Evaluating Fidelity in Prolonged Exposure Dialogues

### v1

@@ Abstract
The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and the high cost of annotation. This work explores the use of synthetic Prolonged Exposure (PE) therapeutic conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models. We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity. We also introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency. Our findings show that although synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it can struggle to capture the subtle dynamics of therapeutic interactions. In our dataset, synthetic dialogues match structural features of real-world dialogues (e.g., speaker switch ratio: 0.98 vs. 0.99), however synthetic interactions do not adequately reflect key fidelity markers (e.g., distress monitoring). We highlight gaps in existing evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures. Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
@@
Suhas BN1  Dominik Mattioli1  Saeed Abdullah1 Rosa I. Arriaga2  Chris W. Wiese3  Andrew M. Sherrill4 1College of Information Sciences and Technology, Penn State University, USA 2School of Interactive Computing, Georgia Tech, USA 3School of Psychology, Georgia Tech, USA 4Department of Psychiatry and Behavioral Sciences, Emory University, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
@@ 1 Introduction
Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022). Real clinical conversations, particularly therapist-client dialogues, are rich sources of data for training models in areas such as mental health diagnostics, therapeutic interventions, and adherence monitoring. However, the high cost of data annotation, privacy concerns surrounding personally identifiable information (PII) (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints in sharing sensitive interactions severely limit access to these datasets. Synthetic data has emerged as a promising alternative to address these challenges (Aher et al., 2023). By simulating conversations in a controlled environment, synthetic data offers scalability and privacy preservation while reducing the reliance on costly real-world annotations.
@@
In trauma-focused mental health care, especially Prolonged Exposure (PE) therapy for PTSD, large language models offer the potential to generate synthetic therapy dialogues at scale. Despite its potential, questions remain about the fidelity and utility of synthetic data to capture the subtle dynamics of therapeutic interactions (e.g., emotional depth, turn-taking patterns, and building therapeutic alliances) (Shen et al., 2024). Recent work (e.g., BN et al. (2025)) has shown that synthetic dialogues can fluently simulate PE sessions across diverse contexts. However, such outputs may only mimic surface structure and tone. Subtle fidelity lapses—like premature therapist reflection or avoidance reinforcement—often remain undetected by automatic metrics or non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig. 1). Without clinical expertise, both humans and automated metrics tend to overestimate fidelity. This paper introduces a fidelity-focused lens to evaluate whether synthetic PE sessions truly adhere to the therapeutic process, not just structure or tone. This fidelity-focused lens involves evaluating dialogues not only on linguistic coherence but also on adherence to specific therapeutic protocols and the successful navigation of key clinical interaction patterns (e.g., managing avoidance, appropriate SUDS monitoring) using both automated metrics and expert clinical assessment.
In this study, we present a systematic comparison of real and synthetic clinical conversations, focusing on their linguistic, structural, and statistical characteristics. Our analysis explores whether synthetic data can replicate key features of real dialogues, including turn-taking patterns, protocol adherence, and emotional progression, while mitigating the risks associated with privacy and data scarcity. By analyzing both the strengths and limitations of synthetic therapy dialogues, our findings inform future improvements in the generation and evaluation of synthetic data for clinical applications. This study contributes to ongoing efforts to bridge the gap between real and synthetic data, providing a roadmap for the future of synthetic dialogue in sensitive domains. To our knowledge, this is the first study to combine statistical, linguistic, and clinical fidelity evaluations on both real and synthetic Prolonged Exposure (PE) therapy conversations at this scale.
@@ 2 Motivation
The development of synthetic clinical dialogues is motivated by several critical challenges in mental healthcare. Prolonged Exposure (PE) therapy—an evidence-based treatment for post-traumatic stress disorder (PTSD)—offers a structured therapeutic approach that could benefit from AI support, but faces significant obstacles:
  1. 1.
    Data scarcity: The limited availability of annotated clinical conversations restricts the development of AI systems for therapist training. Manual annotation is costly and time-consuming, often requiring up to a week per session per annotator, further compounding the scarcity of usable data.
  2. 2.
    Privacy constraints: Patient narratives contain highly sensitive personal information, making clinical data difficult to share or publish under privacy regulations and ethical guidelines.
  3. 3.
    Lack of diversity: Existing datasets often fail to capture diverse demographic profiles, trauma types, and therapeutic approaches, limiting model generalization.
  4. 4.
    Standardization challenges: In the absence of standardized benchmarks, fidelity evaluation of synthetic dialogues remains subjective and inconsistent.
Our work addresses these challenges by creating and systematically evaluating synthetic PE therapy conversations that maintain clinical fidelity while protecting patient privacy. Beyond privacy preservation, synthetic data enables scalable annotation, broader demographic and trauma-type coverage, and controlled variation in dialogue structure—capabilities difficult to achieve with curated real-world data alone. Our framework compares real and synthetic dialogues across linguistic, structural, and clinical fidelity dimensions.
@@ 3 Related Work
The evaluation of synthetic clinical conversation datasets, particularly for psychotherapy like Cognitive Behavioral Therapy (CBT) and general counseling, has grown significantly. However, research focusing on Prolonged Exposure (PE) therapy and other structured trauma-focused treatments remains limited. Existing studies primarily address (1) synthetic dataset generation and fidelity assessment, (2) evaluation metrics beyond lexical similarity, and (3) human-in-the-loop validation, often lacking PE-specific considerations.
@@ 3.1 Synthetic Dataset Generation and Evaluation
Several studies explore generating and assessing synthetic clinical datasets. For instance, Chiu et al. (2024) introduced BOLT to evaluate LLM-generated therapist behaviors (reflections, psychoeducation) in general counseling, while Qiu and Lan (2024) proposed SimPsyDial, benchmarking it against RealPsyDial using the Working Alliance Inventory (WAI-O-S). Zhang et al. (Zhang et al., 2024) presented CPsyCoun, reconstructing dialogues from reports and evaluating them with BERTScore (Zhang et al., 2019), GPTScore, and qualitative review. Others have focused on tailoring dialogues to specific counselor styles (Xie et al., 2024) or enhancing realism via data augmentation (Kim et al., 2024). Although datasets like Thousand Voices simulate structural variation (BN et al., 2025), current evaluations primarily compare synthetic data to real-world general counseling or CBT (Lee et al., 2024a; Shen et al., 2024). They generally lack explicit integration and assessment of trauma-specific markers or the structural fidelity crucial for PE, such as handling avoidance behaviors and managing imaginal exposure sequences. Our work extends these efforts by specifically evaluating whether synthetic PE sessions maintain clinical realism according to PE protocols (e.g., therapist containment, SUDS handling, avoidance redirection), highlighting fidelity violations where linguistic fluency masks procedural errors often missed by standard LLM evaluations or non-clinical reviewers.
@@ 3.2 Beyond Lexical Metrics
Evaluation is shifting from traditional lexical metrics (e.g., BLEU (Papineni et al., 2002)) towards more context-aware approaches. Transformer-based scores like BERTScore and GPTScore assess semantic alignment (Zhang et al., 2024; Xie et al., 2024), and psychological metrics (e.g., CTRS, PANAS in COUNSELINGEVAL (Lee et al., 2024a)) evaluate aspects like empathy and coherence. Specific interaction patterns, like Proactive Questioning Ability (PQA), have also been analyzed (Shen et al., 2024). However, these metrics are often unadapted for trauma-focused contexts, and advanced sequential techniques like Dynamic Time Warping (DTW), potentially useful for assessing turn-by-turn coherence in PE’s structured flow, remain largely unexplored in this domain. While qualitative assessments consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and methods for structured questioning exist (Ren et al., 2024), robust quantitative frameworks for evaluating adherence to PE’s specific therapeutic structure are underdeveloped.
@@ 3.3 Human-in-the-Loop Validation and Engagement Metrics
Human expert validation is recognized as critical. Approaches like COUNSELINGEVAL (Lee et al., 2024a), PsyDT (Xie et al., 2024), and CPsyCoun (Zhang et al., 2024) incorporate clinician or professional reviewer feedback to assess quality, empathy, and safety. Concurrently, personalization and engagement metrics are emphasized (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and prior work addresses empathy (Morris et al., 2018) and structured workflows (Ren et al., 2024). However, these human evaluations predominantly focus on general counseling or CBT contexts. They typically omit systematic assessment of PE-specific elements like trauma cue processing, therapeutic handling of emotional regulation, and adherence to PE protocols, especially when compared against real-world PE therapy data.
@@ 3.4 Bridging the Gap in PE Therapy Evaluation
Prolonged Exposure (PE) therapy is a first-line, empirically supported treatment for PTSD (Sherrill and Rauch, 2019). Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), a significant gap persists in assessing the fidelity of synthetic data for structured, trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024). As discussed, current approaches generally lack the specific focus and metrics needed to evaluate PE’s unique conversational structures (e.g., avoidance handling, trauma cue processing, imaginal exposure flow) and robustness to therapy-specific challenges like patient emotional dysregulation or engagement difficulties (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024). Our work directly addresses this limitation by systematically evaluating synthetic data alignment with structured PE interactions, identifying key areas for improvement in generating clinically valid PE therapy dialogues (See Table 1 and Appendix A).
@@ 4.1.1 System-Level Metrics Comparison
    Linguistic complexity and coherence: Average Utterance Length, Utterance Length Std Dev, and Readability Score assess how natural and readable the synthetic text is. A significant deviation from real data would indicate a lack of coherence in generated responses.
@@ 4.1.2 Correlation Analysis
While mean comparisons provide a general overview, correlation analysis (see Table 3) quantifies the consistency of relationships across metrics between real and synthetic data. This step helps us determine whether patterns observed in real data persist in synthetic dialogues. For instance:
@@ 4.1.3 Statistical Significance Testing
To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 4). This analysis assesses whether observed variations are due to randomness or inherent inconsistencies in the synthetic data. Key takeaways include:
@@ 4.1.4 PE-Specific Metrics
Prolonged Exposure (PE) therapy reduces pathological fear through repeated trauma exposure. To assess the fidelity of synthetic PE sessions, we evaluate key therapeutic constructs (See Appendix A for definitions and Table 5): Trauma Narrative Coherence measures how structured and detailed a client’s trauma account is, reflecting cognitive integration. Emotional Engagement captures the level of emotional expression, linked to better outcomes. Avoidance Handling evaluates how well avoidance behaviors are addressed. Exposure Guidance assesses the therapist’s role in structuring effective exposure exercises. Cognitive Restructuring tracks how clients challenge maladaptive beliefs. Emotional Habituation and SUDS Progression measure distress reduction over repeated exposures. Avoidance Reduction quantifies improvements in engaging with trauma-related content. Emotion Intensity assesses the variability and magnitude of emotional responses. These metrics are derived using linguistic analysis, semantic modeling, and interaction patterns.
@@ 5.1 System-Level Metrics & Correlation
The system-level metrics comparison highlights alignment between real and synthetic dialogues, revealing both structural matches and model-driven differences. While turn-taking patterns (e.g., Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Therapist/Client Turns) remain similar, variations in utterance length, lexical diversity, and entropy-based measures arise due to prolonged exposure therapy and model constraints.
@@
5.2 Statistical Significance Testing
  1. 1.
  2. 2.
  3. 3.
5.3 Feature Importance
Feature importance analysis highlights which metrics most influence synthetic data generation, reinforcing trends observed in prior evaluations. While not the primary focus, these insights provide valuable confirmation of the model’s alignment with conversational structure and linguistic patterns.
  1. 1.
    Turn-taking features contribute heavily: Metrics such as Normalized Speaker Switches and Therapist-Client Turn Ratio emerge as highly influential. This is a positive indication that the model effectively captures conversational flow, ensuring balanced exchanges between therapist and client. The strong correlation between real and synthetic data in these aspects suggests that the generated dialogues maintain natural interaction patterns.
  2. 2.
    Utterance length and entropy-based measures shape responses: Metrics like Average Utterance Length, Perplexity, and Flow Entropy stand out as key distinguishing factors. This aligns with statistical significance results, reflecting that synthetic responses are more structured and predictable. This behavior is expected, as the model prioritizes coherence and consistency, which helps maintain clarity even in complex dialogues. While synthetic responses may be more concise, they still reflect key interaction patterns but require further refinement to support emotionally dynamic exchanges.
  3. 3.
    Lexical richness plays a secondary role: Vocabulary Richness was moderately influential, indicating that synthetic data maintains a varied vocabulary while ensuring structured dialogue. While real-world conversations may exhibit greater linguistic flexibility, synthetic dialogues remain clear and contextually appropriate, balancing variation with coherence.
5.4 Prolonged Exposure Specific Metrics
Table 5 presents the statistical test results comparing synthetic Prolonged Exposure (PE) data with real therapy session data. The results indicate that the synthetic data effectively captures key therapeutic constructs, with several core PE metrics such as Trauma Narrative Coherence, Emotional Engagement, Avoidance Handling, Exposure Guidance, Cognitive Restructuring, Avoidance Reduction, and Emotion Intensity showing strong alignment with real data (p < 0.001) showing strong alignment with real data. These results suggest that the synthetic data successfully replicate the structure and engagement patterns seen in real therapy sessions. However, certain metrics do not reach statistical significance, indicating areas where synthetic data generation may need refinement. Emotional Habituation ($p=0.102$) and SUDS Progression ($p=0.073$) track distress reduction over repeated exposures, and their lack of significance suggests the synthetic data may not fully model this progressive decrease. Similarly, while Trauma Narrative Coherence is well-captured, Narrative Development ($p=0.251$) appears to be less robust, possibly due to limited context retention or rigid structure in generated responses. Despite these discrepancies, the overall performance of synthetic data suggests it is a viable alternative for privacy-sensitive applications. Addressing dynamic narrative evolution could further improve alignment with real data. However, fidelity violations in synthetic sessions are often subtle and may not disrupt structural metrics. For example, therapist utterances can appear empathetic or affirming while inadvertently shifting the session away from trauma anchoring. These moments often go unnoticed by automated scoring or non-clinical reviewers, underscoring the need for fidelity-aware evaluation frameworks that integrate human clinical judgment.
6 Discussion
In this paper, we evaluate the fidelity of synthetic Prolonged Exposure (PE) therapy conversations by comparing them to real interactions using linguistic, structural, and statistical analyses. We discuss the findings from four key perspectives: system-level metrics and correlation (Table 2), statistical significance testing (Table 3), feature importance (Table 4), and PE-specific metrics (Table 5).
6.1 Clinical Implications
Our findings have several important implications for clinical practice. First, synthetic PE sessions could serve as valuable training tools for novice therapists learning to identify protocol deviations, as they provide controlled examples of both adherent and non-adherent interactions without privacy concerns. Second, the consistent replication of structural elements (turn-taking, session flow) suggests synthetic data can supplement clinical training materials, particularly in regions with limited access to specialized trauma training. However, the identified limitations in capturing emotional habituation patterns and dynamic narrative development indicate synthetic conversations should not replace supervised clinical practice with real clients. The subtle fidelity violations we identified (Figure 1) would likely go unnoticed by non-specialist reviewers, highlighting the need for expert-guided evaluation of AI tools in clinical settings. Finally, our PE-specific metrics provide a framework for clinicians and developers to systematically evaluate AI-generated content for trauma treatment, potentially establishing minimum standards for therapeutic applications of synthetic data.
6.2 Supplementing Automated Metrics with Qualitative Fidelity Analysis
To supplement automated metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using Prolonged Exposure (PE) clinical guidelines. This qualitative analysis is essential, as statistical metrics alone may not detect subtle fidelity violations present even in linguistically fluent outputs. Fig. 1 presents five representative examples highlighting moments where such fluent outputs diverge from therapeutic fidelity, potentially undermining therapeutic efficacy. These include instances of role drift, premature processing, failure to maintain focus on exposure targets, improper SUDS implementation, and missed containment opportunities—violations that would often be imperceptible to both LLMs and lay reviewers. Fig. 1 illustrates these subtle violations clearly:
  • In Ex. 1, we observe role drift, where the therapist assumes the client’s narrative responsibility. While maintaining conversational coherence, this action undermines the therapeutic goal of client-driven processing.
  • Ex. 2 demonstrates a failure to maintain focus on exposure targets, allowing avoidance behavior that linguistic metrics cannot distinguish from therapeutically appropriate responses.
  • Ex. 3 and 4 show premature reflection and intervention, where the therapist interjects before full emotional processing occurs. This pattern can appear empathetic but critically disrupts the habituation processes central to PE efficacy.
  • Ex. 5 highlights improper SUDS implementation, which maintains the structural dialogue flow but compromises the validity of the clinical assessment.
These specific examples underscore why evaluation frameworks must integrate clinical expertise alongside computational metrics to accurately assess therapeutic fidelity in synthetic PE sessions.
7 Future Work
8 Conclusions
9 Ethical Concerns
Limitations
Appendix A Metric Definitions
We summarize some common PE terms  (Foa et al., 2007; Powers et al., 2010; Rauch et al., 2009; Hembree et al., 2003) below:
  1. 1.
    SUDS Progression: Change in Subjective Units of Distress Scale (SUDS) values mentioned by the client across the session.
  2. 2.
    Emotional Habituation: A decrease in distress levels or emotional intensity from the beginning to end of imaginal exposure.
  3. 3.
    Trauma Narrative Coherence: The syntactic and semantic coherence of the client’s trauma story, as measured by discourse coherence metrics.
  4. 4.
    Emotional Engagement: Captures the level of emotional expression, linked to better outcomes.
  5. 5.
    Avoidance Handling: Evaluates how well avoidance behaviors are addressed.
  6. 6.
    Exposure Guidance: Assesses the therapist’s role in structuring effective exposure exercises.
  7. 7.
    Cognitive Restructuring: Tracks how clients challenge maladaptive beliefs.
  8. 8.
    Avoidance Reduction: Quantifies improvements in engaging with trauma-related content.
  9. 9.
    Emotion Intensity: Assesses the variability and magnitude of emotional responses.
Appendix B Additional Tables

### v3

@@ Abstract
The growing adoption of synthetic data in healthcare is driven by privacy concerns, limited access to real-world data, and high annotation costs. This work explores the use of synthetic Prolonged Exposure (PE) therapy conversations for Post-Traumatic Stress Disorder (PTSD) as a scalable alternative for training and evaluating clinical models. We systematically compare real and synthetic dialogues using linguistic, structural, and protocol-specific metrics, including turn-taking patterns and treatment fidelity. We introduce and evaluate PE-specific metrics derived from linguistic analysis and semantic modeling, offering a novel framework for assessing clinical fidelity beyond surface fluency. Our findings show that while synthetic data holds promise for mitigating data scarcity and protecting patient privacy, it often struggles to capture the subtle dynamics of therapeutic interactions. Synthetic therapy dialogues closely match the structural features of real conversations (e.g., speaker switch ratio: 0.98 vs. 0.99), but often fails to adequately reflect key fidelity markers such as distress monitoring. This work highlights gaps in current evaluation frameworks and advocate for fidelity-aware metrics that go beyond surface fluency to uncover clinically significant failures. Our findings clarify where synthetic data can effectively complement real-world datasets—and where critical limitations remain.
@@
Suhas BN1  Dominik Mattioli1  Andrew M. Sherrill2 Rosa I. Arriaga3  Chris W. Wiese4  Saeed Abdullah1 1College of Information Sciences and Technology, Penn State University, USA 2Department of Psychiatry and Behavioral Sciences, Emory University, USA 3School of Interactive Computing, Georgia Tech, USA 4School of Psychology, Georgia Tech, USA {bnsuhas,saeed}@psu.edu, andrew.m.sherrill@emory.edu
@@ 1 Introduction
Training machine learning models in sensitive domains like healthcare remains a challenge (Giuffrè and Shung, 2023; Kokosi and Harron, 2022). Access to real clinical conversations—crucial for modeling tasks like mental health diagnostics and therapeutic dialogue understanding—is severely limited by high annotation costs, patient privacy concerns (Kokosi and Harron, 2022; BN and Abdullah, 2022; BN et al., 2023), and ethical constraints on data sharing. Synthetic data has emerged as a promising alternative, offering scalability and privacy preservation while reducing dependence on real-world annotations (Aher et al., 2023).
@@
In trauma-focused mental health care, particularly Prolonged Exposure (PE) therapy for PTSD, large language models (LLMs) can generate synthetic therapy dialogues at scale. However, questions remain about whether these dialogues capture more than surface-level fluency—specifically, the subtle dynamics of therapeutic fidelity such as emotional pacing, avoidance management, and protocol adherence (Shen et al., 2024). To address this, we develop and validate methods to measure clinically-relevant fidelity in generated dialogues—moving beyond standard metrics like coherence or perplexity. Prior work shows that while synthetic PE sessions can convincingly mimic real sessions in tone and structure (BN et al., 2025), they may still commit fidelity lapses, such as premature reflection or reinforcement of avoidance, which often go unnoticed by both automatic metrics and non-clinical annotators (Chiu et al., 2024; Zhang et al., 2024; Lee et al., 2024b).
For instance, a therapist saying “That’s a really powerful insight” mid-exposure may appear empathic but violates PE protocol by derailing trauma processing (see Fig. 1). Without clinical expertise, both humans and automated metrics tend to overestimate fidelity. We introduce a fidelity-focused lens that evaluates dialogues on multiple dimensions: linguistic coherence, adherence to PE protocols, and the therapist’s navigation of key clinical interactions (e.g., managing avoidance, SUDS monitoring). This framework integrates automated scoring and expert clinical assessment for a more rigorous evaluation of synthetic dialogue quality. We present the first large-scale, multi-dimensional comparison of real and synthetic PE therapy sessions—analyzing their linguistic, structural, and clinical fidelity characteristics. Our findings demonstrate the current strengths and limitations of synthetic dialogues and inform future improvements in generation and evaluation methods. Our study contributes to bridging the real–synthetic data gap and provides a roadmap for advancing synthetic dialogue in sensitive healthcare domains.
@@ 2 Motivation
Developing synthetic clinical dialogues is essential to overcome persistent barriers in mental health AI. In the context of Prolonged Exposure (PE) therapy, four challenges stand out: (1) data scarcity, due to the high cost and effort of annotation; (2) privacy constraints, which limit access to sensitive patient narratives; (3) lack of diversity, with datasets often failing to capture varied trauma types and demographics; and (4) evaluation inconsistency, as fidelity assessment lacks standardized benchmarks.
This work directly addresses these issues by generating and evaluating synthetic PE dialogues that preserve protocol fidelity, enable scalable annotation, and support more equitable and robust training data for mental health applications.
@@ 3 Related Work
While synthetic clinical conversation datasets have advanced evaluations for general counseling and CBT, work specific to Prolonged Exposure (PE) and structured trauma-focused therapy is limited. Most studies focus on (1) synthetic dataset generation, (2) evaluation metrics beyond lexical similarity, and (3) human-in-the-loop validation, but generally lack PE-specific considerations.
@@ 3.1 Synthetic Dataset Generation and Evaluation
Recent work has explored synthetic clinical dialogue generation and assessment. BOLT evaluates LLM-generated therapist behaviors in general counseling (Chiu et al., 2024), while SimPsyDial benchmarks synthetic data using the Working Alliance Inventory (Qiu and Lan, 2024). CPsyCoun reconstructs dialogues for evaluation with BERTScore, GPTScore, and qualitative review (Zhang et al., 2024, 2019). Other studies focus on counselor style (Xie et al., 2024) and data augmentation (Kim et al., 2024). While Thousand Voices of Trauma introduces structural variation (BN et al., 2025), most evaluations compare to general counseling or CBT (Lee et al., 2024a; Shen et al., 2024) and rarely assess trauma-specific markers or PE’s structural fidelity, such as avoidance handling and imaginal exposure sequences. Our work directly addresses this by evaluating whether synthetic PE sessions capture clinical realism per PE protocols (e.g., SUDS, avoidance redirection), highlighting procedural errors missed by standard metrics.
@@ 3.2 Beyond Lexical Metrics
Evaluation is shifting from lexical metrics like BLEU (Papineni et al., 2002) to semantic ones such as BERTScore and GPTScore (Zhang et al., 2024; Xie et al., 2024). Psychological measures (CTRS, PANAS) in COUNSELINGEVAL assess empathy and coherence (Lee et al., 2024a), and interactional features like PQA are explored (Shen et al., 2024). However, most metrics are not adapted for trauma therapy, and methods like Dynamic Time Warping (DTW)—potentially useful for structured PE flows—remain underused. While some qualitative reviews consider emotional tone and goal alignment (Zhang et al., 2024; Xie et al., 2024), and structured questioning is explored (Ren et al., 2024), robust frameworks for quantitative evaluation of PE-specific structure are lacking.
@@ 3.3 Human-in-the-Loop Validation and Engagement Metrics
Clinician validation is now common, as in COUNSELINGEVAL (Lee et al., 2024a), PsyDT (Xie et al., 2024), and CPsyCoun (Zhang et al., 2024), which assess quality, empathy, and safety. Metrics for engagement and personalization are also used (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024). Prior work covers empathy (Morris et al., 2018) and structured workflows (Ren et al., 2024). However, human evaluation remains focused on general counseling or CBT, rarely addressing PE-specific components like trauma cue processing or protocol adherence compared to real PE data.
@@ 3.4 Bridging the Gap in PE Therapy Evaluation
PE is a first-line treatment for PTSD (Sherrill and Rauch, 2019; Varkovitzky et al., 2018; Rauch et al., 2021; Ragsdale et al., 2020; Evans et al., 2020; Yasinski et al., 2017). Despite progress in synthetic dialogue generation (Qiu and Lan, 2024; Zhang et al., 2024; Xie et al., 2024), evaluation metrics (Lee et al., 2024a; Zhang et al., 2024; Xie et al., 2024), and human validation, a gap remains in fidelity assessment for trauma-focused therapies like PE (Chiu et al., 2024; Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Hu et al., 2024). Current approaches lack metrics and focus tailored to PE’s unique conversational structure—including avoidance handling, trauma cue processing, and imaginal exposure flow—and do not address challenges such as patient emotional dysregulation (Qiu and Lan, 2024; Lee et al., 2024a; Shen et al., 2024; Zhang et al., 2024; Xie et al., 2024; Kim et al., 2024). Our work systematically evaluates synthetic data for alignment with PE structure and clinical validity, identifying areas for improving synthetic PE dialogue quality (see Table 1, Appendix A).
@@ 4.1.1 System-Level Metrics Comparison
    Linguistic complexity and coherence: Average Utterance Length, Utterance Length Std Dev, and Readability Score assess how natural and readable the synthetic text is. Significant deviations indicate poor coherence.
@@ 4.1.2 Correlation Analysis
While mean comparisons provide a general overview, correlation analysis (see Table 3) quantifies the consistency of relationships across metrics between real and synthetic data. This determines if real-data patterns persist in synthetic dialogues. For instance:
@@ 4.1.3 Statistical Significance Testing
To confirm whether differences between real and synthetic data are statistically meaningful, we conduct Mann-Whitney U tests (Table 4). This assesses if variations are random or due to inherent inconsistencies. Key takeaways include:
@@ 4.1.4 PE-Specific Metrics
PE therapy reduces pathological fear through repeated trauma exposure. To assess the fidelity of synthetic PE sessions, we evaluate key therapeutic constructs (See Appendix A for definitions and Table 5): Trauma Narrative Coherence measures how structured and detailed a client’s trauma account is, reflecting cognitive integration. Emotional Engagement captures the level of emotional expression, linked to better outcomes. Avoidance Handling evaluates how well avoidance behaviors are addressed. Exposure Guidance assesses the therapist’s role in structuring effective exposure exercises. Cognitive Restructuring tracks how clients challenge maladaptive beliefs. Emotional Habituation and SUDS Progression measure distress reduction over repeated exposures. Avoidance Reduction quantifies improvements in engaging with trauma-related content. Emotion Intensity assesses the variability and magnitude of emotional responses. These metrics are derived using linguistic analysis, semantic modeling, and interaction patterns.
4.1.5 Qualitative Fidelity Assessment
To supplement automated evaluation metrics, we conducted a manual review of synthetic dialogues, annotating exchanges for fidelity adherence or violations using established PE clinical guidelines. Manual fidelity annotations were then reviewed by a licensed clinical psychotherapist for adherence to PE protocol. While inter-rater agreement was not computed in this exploratory phase, these annotations served as a qualitative tool to identify and illustrate typical fidelity lapses—especially those potentially overlooked by automated metrics, rather than as quantitative ground truth. Representative examples of such lapses identified through this process are illustrated in Figure 1 and discussed in Section 6.2.
@@ 5.1 System-Level Metrics & Correlation
The system-level metrics comparison highlights alignment between real and synthetic dialogues, revealing both structural matches and model-driven differences. While turn-taking patterns (e.g., Normalized Speaker Switches, Therapist-Client Turn Ratio, and Normalized Therapist/Client Turns) remain similar, variations in utterance length, lexical diversity, and entropy-based measures arise due to PE therapy and model constraints.